vector r.n=q ,find the intersecting part of x axis.tell me the answer

Grade:12

1 Answers

Aditya Gupta
5 years ago
the above eqn r.n=q seems to be a plane eqn.
let n= (n1,n2,n3)
when the plane intersects x axis then y and z coordinates are zero.
so x0*n1=q
or x0=q/n1 but we know that n1= n.i
so, it intersects x axis at
(q/n.i,0,0)